OpenStudy (anonymous):

The base of a triangle is 4cm greater then the height. The area is 48cm^3. Find the height and the length of the base. The height of the triangle is _cm The base if the triangle is _ cm Please help

OpenStudy (anonymous):

The base a is h+4cm, area is (a h)/2, so ((h+4) h)/2=48, then you get h^2+4h=96, and when you solve quadratic equation by h, you get h=8cm (you should take positive value of h) and a = 8+4=12cm.

OpenStudy (anonymous):

thanks very much :)